the width of a rectangle is 3 feet less than its length. if the perimeter is 22 feet, what are both its length and width?
Found 2 solutions by Mathtut, midwood_trail:
Answer by Mathtut(3670)

P=2(L+W) W=L-3 substituting W and P(given as 22ft) into the first equation we get 22=2(L+(L-3) so 22=2(2L-3) distributing the right side of the equation we arrive at 22=4L-6. Add 6 and divide by 4 on each side of the equation and we get L=7 since L=7 then W=L(7)-3 so W=4
L=7
W=4

Answer by midwood_trail(310)

The width of a rectangle is 3 feet less than its length. if the perimeter is 22 feet, what are both its length and width?
width = x - 3
length = x
perimeter = 22
P = 2L + 2W
22 = 2x + 2(x - 3)
22 = 2x + 2x - 6
22 = 4x - 6
22 + 6 = 4x
28 = 4x
28/4 = x
7 = x
Since x = length, then length is 7 feet.
Our width is x - 3.
Replacing x with 7 and subtracting 3, will give us 4 feet.
Final answer: width = 4 feet and length = 7 feet.